a group of girls share a packet of sweets. if each girl gets 5 sweets, there will be 4 extra sweets. If each girl gets 6 sweets, there will be 5 sweets short. How many girls are there?
Answer by ikleyn(52920)

.
Let N be the number of girls.


From the condition,

5N + 4 = 6N -5      (both sides are the number of sweets calculated different ways).


From the equation,  4 + 5 = 6N - 5N,   or  N = 9.



Answer.  9 girls.